OpenStudy (credmond):

Can someone please help me step by step with this question? I know you have to use Henderson-Hasselbalch's equation (pH = pKa + log [A-] /[HA]) but I don't know where to start/finish this problem. Thanks! :) 100 mL of buffer contains 0.041 moles of HC2H3O2 (acetic acid) and 0.038 moles of NaC2H3O2 (sodium acetate). The Ka of acetic acid is 1.8 x 10-5. Using the above equation, what is the pH of this solution?

OpenStudy (vincent-lyon.fr):

Simply plug in the values given into HH's equation, remembering that pKa = - lg Ka

OpenStudy (aaronq):

Substitute your values in: \(pH=pKa+log\dfrac{[C_2H_3O_2^-] }{[HC_2H_3O_2]} \rightarrow pH=-log(1.8 x 10^{-5})+log\dfrac{(\dfrac{0.038 moles}{0.1L}) }{(\dfrac{0.041 moles}{0.1L})}\) plug it into your calculator

OpenStudy (credmond):

THANK YOU SO MUCH!! :) LIFE SAVER!!